B.WATER SAMPLE ANALYSIS.
Water temperature 27°C.
SAMPLE 1
a)Volume of Na2S2O3 required to titrate water sample; Final buret reading : 12.7mL, Initial buret reading= 2.6mL,Volume of Na2S2O3=10.1mL.
b)Moles of Na2S2O3 required to titrate water sample(show calculations)=_________mol.
c)Moles of O2 present in 200mL of water sample(show calculations)________mol.
d)Grams of O2 present in 200mL of water sample=_______g.
e)Milligrams of O2 present in 200mL of water sample=________mg.
f)Milligrams of O2 present in 1000mL of water sample=______ppm.
g)Saturation level for yourwater sample(show calculations)=_________ppm.
1
Expert's answer
2019-12-18T06:12:50-0500

b) n(Na2S2O3) = C(Na2S2O3) x V = C(Na2S2O3) mol/L x 0.0101 L

c) 1 mol of O2 reacts with 4 mol S4O62- (8 mol S2O32-)

d) m(O2) = n(O2) x Mr(O2) = n(O2) mol x 32 g/mol

e) 1 gram = 1000 mg

m(O2) x 1000

f) In 1000 mL O2 m(O2) x 1000 x 5

g) Percent saturation (% saturation) is the amount of dissolved oxygen in the water sample compared to the maximum amount that could be present at the same temperature